OpenStudy (anonymous):

Find the slope of points R(-2,-3), S(-2,-5)

OpenStudy (anonymous):

(y2-y1)/(x2-x1)=(-3-(-5))/(-2-(-2)=2/0=undefined

OpenStudy (anonymous):

the graph for that would be a vertical line, which indicates that the slope is undefined

OpenStudy (anonymous):

Ohh okay thanks I was confused :P